# View-Centric Multi-Object Tracking with Homographic Matching in UAV Scene

### v2

@@ Abstract
In this paper, we address the challenge of multi-object tracking (MOT) in moving Unmanned Aerial Vehicle (UAV) scenarios, where irregular flight trajectories, such as hovering, turning left/right, and moving up/down, lead to significantly greater complexity compared to fixed-camera MOT. Specifically, changes in the scene background not only render traditional frame-to-frame object IOU association methods ineffective but also introduce significant view shifts in the objects, which complicates tracking. To overcome these issues, we propose a novel universal HomView-MOT framework, which for the first time, harnesses the view Homography inherent in changing scenes to solve MOT challenges in moving environments, incorporating Homographic Matching and View-Centric concepts. We introduce a Fast Homography Estimation (FHE) algorithm for rapid computation of Homography matrices between video frames, enabling object View-Centric ID Learning (VCIL) and leveraging multi-view Homography to learn cross-view ID features. Concurrently, our Homographic Matching Filter (HMF) maps object bounding boxes from different frames onto a common view plane for a more realistic physical IOU association. Extensive experiments have proven that these innovations allow HomView-MOT to achieve state-of-the-art performance on prominent UAV MOT datasets VisDrone and UAVDT.
1 Introduction
Multi-object Tracking(MOT) has been a classical and longstanding task in various wide applications [29, 30]. It is aimed to address the cross-frame trajectory of each object in video frames. Various MOT paradigms have been proposed for conventional static scene captured by fixed cameras [2, 21, 40]. Recently, benefited from the advancement of Unmanned Aerial Vehicle (UAV) technologies, the accessibility and analysis of scenes under moving UAV views is opening new horizons for the community [22, 56].
@@ 1 Introduction
In this paper, we envision utilizing Homography property of spatial scenes to address the aforementioned issues and propose a novel View-Centric Multi-Object Tracking with Homographic Matching (HomView-MOT) in Moving UAV scene. Firstly, for each UAV-captured video sequence, we regard it as multi-view shooting of one or more scenes. Therefore, we initially compute the Homography transformation matrix between adjacent frames. Notably, to enhance tracking efficiency, we avoid calculating between every two adjacent frames and instead perform skip-frame computation, only calculating the Homography matrix for every $h$ frames apart. We design an efficient Fast Homography Estimation (FHE) algorithm to reasonably estimate the Homography matrix for every two frames within this interval. Secondly, based on the inter-frame scene Homography matrix, we present the Homographic Matching Filter (HMF), which can map detected objects within them to the same viewpoint space to calculate the "physical IoU in the real world" of objects between two frames (as shown in Figure 2), thereby improving the precision of track matching. Finally, to address the issue of object view changes during UAV flight, we propose View-Centric ID Learning (VCIL). By combining the inter-frame scene Homography matrix, we design Homographic Slot Attention (HSA) to explicitly extract view-centric slots for each frame and correlate views between adjacent frames, thus updating object ID features to cope with view changes during UAV flight. By integrating these designs, we make HomView-MOT an efficient and effective multi-object tracker.
@@
  • For the first time, we address the problem of multi-object tracking (MOT) in moving scenarios by integrating the concept of Homography from different views, introducing Homographic Matching and View-Centric concepts into the field and proposing a universal HomView-MOT framework.
  • Specifically, we begin by employing a Fast Homography Estimation (FHE) algorithm to rapidly estimate the Homography matrix between any two video frames. Based on this, we perform object View-Centric ID Learning (VCIL), utilizing the Homography of multiple perspectives to learn cross-view ID features. Furthermore, leveraging the Homography matrix, we design a Homographic Matching Filter (HMF) to map the object bounding boxes from different frames onto the same view plane, allowing for more accurate physical IoU associations that closely mirror the real world.
@@
2 Related Work
@@
Multi-Object Tracking. Classical MOT algorithms [2, 34, 4, 9] generally adopt a tracking by detection framework, where the first step is to detect all object boxes in each frame, and the second step involves object ID matching between adjacent frames through data association[42, 34]. In the first detection step, related algorithms such as SORT[2], Deep SORT[34], and others typically use classical detection models like Fast-RCNN[10] and YOLO[26]. Benefiting from the recent progress in deep learning[60, 58, 57, 59, 51, 54, 52, 53, 55, 18, 16, 17, 15, 13, 12, 14, 49, 46, 48, 47, 38, 50, 37, 39, 31, 32], recent research has also introduced vision transformer detectors into MOT to achieve higher performances, as seen in TransTrack[28], TrackFormer[24], and similar approaches. In the second step of data association, traditional MOT algorithms usually rely on Matching Algorithms based on IoU association between target boxes in adjacent frames. This method is quite effective in static scenes captured by fixed cameras. However, in UAV footage and the like, the moving scene background, causing different frames to be on different viewing planes, can render the conventional IoU association method inaccurate. To address this, UAVMOT[22] proposes an Adaptive Motion Filter to judge the different motion states of the UAV and process them differently. In this paper, we propose a Homography Matching Filter that, based on Homography transformation, maps the object boxes from adjacent frames in different perspectives into the same viewpoint space before performing IoU association to enhance the robustness of tracking.
ID Embedding Assistance. To enhance robustness, many MOT frameworks propose adding an embedding based on ReID learning on top of detection, which is used to improve multi-object data association[43, 22]. In scenes captured by stationary cameras, targets generally present a very stable viewpoint, making ID learning comparatively simple and straightforward. However, in the case of UAVs, due to the variability of shooting angles, an object may exhibit a rich array of viewpoints. Therefore, we further propose to explicitly learn the view characteristics of targets to optimize the feature representation of object IDs.
3 Method
3.1 Overview
@@
Figure 3: The overview of our proposed HomView-MOT, adhering to the classic tracking-by-detection paradigm. It begins with a Fast Homography Estimation (FHE) algorithm that rapidly calculates the Homography transformation matrix between any two adjacent frames for view scene projection. With a shared feature extraction network, we obtain the objects detection boxes ($B_{t}$, $B_{t+1}$) and ID features ($ID_{t}$, $ID_{t+1}$). Subsequently, the View-Centric ID Learning (VCIL) is utilized to explicitly assimilate the multi-view information and update the ID features. For tracking, the Homography Matching Filter (HMF) is employed to project the object boxes from adjacent frames onto a common view planar space for IoU association. This is coupled with ID association and processed through Matching Algorithm to produce final tracking results.
3.2 Fast Homography Estimation for View Projection
As described in Sec. 1, when the UAV flight state changes (e.g. hovering, moving up/down, turning left/right) occur, the captured scene background will undergo significant changes, making adjacent frames appear as different view planes over the same scene. To facilitate cross-view learning and IoU association for object IDs, we propose to perform view projection for these two planes, in the form of Homography matrix which is extensively applied to various 3D applications [44]. Generally, the Homography matrix for every two adjacent frames can be calculated with the Homography Estimation (HE) algorithm. For instance, as shown in Figure 4, we use a shared SuperPoint [6] network to extract keypoints from both frames separately. The distribution of these keypoints is able to represent the view information of the scene [41, 20, 45, 19]. Then, we can estimate the Homography matrix based on the matching of keypoints between the two frames by the SuperGlue algorithm [27].
@@ 3.2 Fast Homography Estimation for View Projection
However, directly performing Homography estimation calculation for every two frames is very time-consuming and would greatly reduce tracking efficiency. Therefore, we propose a Fast Homography Estimation (FHE) method, where we only need to directly calculate the Homography matrices for a limited number of key frames, while the Homography matrices for other adjacent frames can be fast estimated using the ones of the key frames. Specifically, we utilize a uniform sampling process to select the key frames from the input video sequence $I_{t}(t\in[1,T])$, the sampling interval is set to $h$. We only perform a direct calculation of the Homography matrices $\{H_{t,t+h},H_{t+h,t+2h},...\}$ between the sampled frames, where $H_{t,t+h}$ indicates the Homography matrix from the $t+h$-th to $t$-th frame. For any two frames , based on the Derivation 1 in the appendix, we prove that their Homography matrix can be reasonably estimated through the derivation of Homography matrix of their adjacent key frames.
Specifically, as shown in Figure 5, We select three highly representative frames $I_{t_{1}},I_{t_{2}},I_{t_{3}}$ to illustrate the computation process of the Homography matrix between any two frames, where $k_{1}<t_{1}<t_{2}<k_{2}<t_{3}<k_{3}$, and $k_{1},k_{2},k_{3}$ are the sampled frames. The core calculations are as follows:
@@
    • From a non-sampled frame to any frame in the previous sampling interval: $H_{t_{2},t_{1}}=H_{t_{2},k_{2}}H_{k_{2},t_{1}}$ and $H_{t_{2},k_{1}}=H_{t_{2},k_{2}}H_{k_{2},k_{1}}$. The calculation of $H_{t_{2},k_{2}}$ is similar to the previous criterion.
The aforementioned computation process involves only interpolation calculations and matrix multiplication, which ensures high computational efficiency in practical implementation. Through the derivations described above, we can rapidly and cost-effectively calculate the Homography matrix between any two frames.
3.3 View-Centric ID Learning
@@
${\rm HSA}_{t}$ $={\rm Softmax}(\frac{1}{\sqrt{C}}k(F_{t}^{ID})\times(H_{t,t+1}\cdot q(S_{t})^{\top}))\times v(F_{t}^{ID})$ (1)
${\rm HSA}_{t+1}$ $={\rm Softmax}(\frac{1}{\sqrt{C}}k(F_{t+1}^{ID})\times(H_{t+1,t}\cdot q(S_{t+1})^{\top}))\times v(F_{t+1}^{ID})$
where $q(\cdot),k(\cdot),v(\cdot)$ are the linear projections. $F_{t}^{ID},F_{t+1}^{ID}$ are the object ID features in frame $t,t+1$, respectively. $S_{t},S_{t+1}$ are the View-Centric Slots in frame $t,t+1$, respectively. $\top$ is the matrix transposition operator. $C$ is the feature dimension. Then we correlate the cross-frame views by,
$S^{\prime}={\rm Softmax}(\frac{S_{t}\times S_{t+1}^{\top}}{\sqrt{C}})\times S_{t+1}$ (2)
@@
3.4 Homography Matching Filter

### v3

# View-Centric Multi-Object Tracking with Homographic Matching in Moving UAV

@@ Abstract
In this paper, we address the challenge of Multi-Object Tracking (MOT) in moving Unmanned Aerial Vehicle (UAV) scenarios, where irregular flight trajectories, such as hovering, turning left/right, and moving up/down, lead to significantly greater complexity compared to fixed-camera MOT. Specifically, changes in the scene background not only render traditional frame-to-frame object IoU association methods ineffective but also introduce significant view shifts in the objects, which complicates tracking. To overcome these issues, we propose a novel HomView-MOT framework, which for the first time, harnesses the view homography inherent in changing scenes to solve MOT challenges in moving environments, incorporating homographic matching and view-centric concepts. We introduce a Fast Homography Estimation (FHE) algorithm for rapid computation of homography matrices between video frames, enabling object View-Centric ID Learning (VCIL) and leveraging multi-view homography to learn cross-view ID features. Concurrently, our Homographic Matching Filter (HMF) maps object bounding boxes from different frames onto a common view plane for a more realistic physical IoU association. Extensive experiments have proven that these innovations allow HomView-MOT to achieve state-of-the-art performance on prominent UAV MOT datasets VisDrone and UAVDT.
I Introduction
Multi-Object Tracking (MOT) has been a classical and longstanding task in various wide applications [1, 2]. It is aimed to address the cross-frame trajectory of each object in video frames. Various MOT paradigms have been proposed for conventional static scene captured by fixed cameras [3, 4, 5]. Recently, benefited from the advancement of Unmanned Aerial Vehicle (UAV) technologies [6, 7, 8], the accessibility and analysis of scenes under moving UAV views is opening new horizons for the community [7, 9].
@@ I Introduction
In this paper, we envision utilizing homography property of spatial scenes to address the aforementioned issues and propose a novel View-Centric Multi-Object Tracking with Homographic Matching (HomView-MOT) in Moving UAV scene. Firstly, for each UAV-captured video sequence, we regard it as multi-view shooting of one or more scenes. Therefore, we initially compute the homography transformation matrix between adjacent frames. Notably, to enhance tracking efficiency, we avoid calculating between every two adjacent frames and instead perform skip-frame computation, only calculating the homography matrix for every $h$ frames apart. We design an efficient Fast homography Estimation (FHE) algorithm to reasonably estimate the homography matrix for every two key frames within this interval. Secondly, based on the inter-frame scene homography matrix, we present the Homographic Matching Filter (HMF), which can map detected objects within them to the same viewpoint space to calculate the "physical IoU in the real world" of objects between two frames (as shown in Figure 1), thereby improving the precision of track matching. Finally, to address the issue of object view changes during UAV flight, we propose View-Centric ID Learning (VCIL). By combining the inter-frame scene homography matrix, we design Homographic Slot Attention (HSA) to explicitly extract view-centric slots for each frame and correlate views between adjacent frames, thus updating object ID features to cope with view changes during UAV flight. By integrating these designs, we make HomView-MOT an efficient and effective multi-object tracker.
@@
  • For the first time, we address the problem of Multi-Object Tracking (MOT) in moving UAV scenarios by integrating the concept of homography from different views, introducing Homographic Matching and View-Centric concepts into the field and proposing a universal HomView-MOT framework.
  • Specifically, we begin by employing a Fast Homography Estimation (FHE) algorithm to rapidly estimate the homography matrix between any two key frames. Based on this, we perform object View-Centric ID Learning (VCIL), utilizing the homography of multiple perspectives to learn cross-view ID features. Furthermore, leveraging the homography matrix, we design a Homographic Matching Filter (HMF) to map the object bounding boxes from different frames onto the same view plane, allowing for more accurate physical IoU associations that closely mirror the real world.
@@
Figure 3: The overview of our proposed HomView-MOT. It begins with a Fast Homography Estimation (FHE) algorithm that rapidly calculates the homography transformation matrix between any two adjacent frames for view scene projection. Subsequently, the View-Centric ID Learning (VCIL) is utilized to explicitly assimilate the multi-view information and update the ID features. For tracking, HMF is employed to project the object boxes from adjacent frames onto a common view planar space for IoU association. This is coupled with ID association and processed through Matching Algorithm to produce final tracking results.
II Related Work
II-A Multi-Object Tracking
@@
Classical MOT algorithms generally adopt a tracking by detection framework [3, 11, 12, 13, 14, 15], where the first step is to detect all object boxes in each frame in videos[16, 17, 18], and the second step involves object ID matching between adjacent frames through data association[19, 11]. In the first detection step, related algorithms such as SORT[3], Deep SORT[11]. Recent research has also introduced vision transformer detectors into MOT to achieve higher performances, as seen in TransTrack[20] and TrackFormer[21]. In the second step of data association, traditional MOT algorithms usually rely on Matching Algorithms based on IoU association between target boxes in adjacent frames. This method is quite effective in static scenes captured by fixed cameras. However, in UAV footage and the like, the moving scene background, causing different frames to be on different viewing planes [6, 22, 8], can render the conventional IoU association method inaccurate. To address this, UAVMOT[7] proposes an Adaptive Motion Filter to judge the different motion states of the UAV and process them differently. Moreover, ASTCA [23] incorporates spatial-temporal context to mitigate background interference and the boundary effect in UAV tracking, which is a challenge also pertinent to MOT in complex aerial views. UDALT [24] offers a fascinating perspective on reducing annotation costs by selecting the most informative samples. Self-SDCT [25] utilize multi-cycle consistency loss as self-supervised signal to learn feature representations from adjacent frames without manual annotations is highly relevant.
II-B ID Embedding Assistance and View-Oriented Learning
To enhance robustness, many Multi-Object Tracking (MOT) frameworks propose adding an embedding based on ReID (re-identification) learning [26, 27, 28, 29, 30, 31, 32] on top of detection, which is used to improve multi-object data association [33, 34, 7] in videos [35, 36, 37]. For instance, FairMOT [33] integrates joint detection and ReID learning into a unified framework, achieving state-of-the-art performance in pedestrian tracking tasks. Similarly, UAVMOT [7] focuses on addressing challenges specific to aerial views, such as scale variations and occlusions, by leveraging robust ReID embeddings. In scenes captured by stationary cameras, targets generally present a very stable viewpoint, making ID learning comparatively simple and straightforward. This stability allows traditional methods like DeepSORT [38] to rely heavily on appearance features for accurate data association. However, in the case of UAVs (Unmanned Aerial Vehicles), due to the variability of shooting angles, an object may exhibit a rich array of viewpoints. This complexity introduces significant challenges for conventional ReID approaches, as highlighted in studies like [39, 40, 41], which emphasize the need for viewpoint-invariant representations in aerial scenarios.
II-C Camera Motion Compensation
The challenge of camera motion in Multi-Object Tracking (MOT) has been actively addressed by Camera Motion Compensation (CMC) techniques. For instance, UCMCTrack [42] introduces a uniform compensation model and a ground plane-based Mapped Mahalanobis Distance, demonstrating robustness in various scenarios but potentially struggling with the non-uniform, irregular motion patterns characteristic of UAV flights. Similarly, another line of research focuses on Global Motion Compensation (GMC) for UAVs [43, 44], often employing frame-by-frame feature matching to align consecutive images and mitigate the impact of camera movement. While these methods effectively stabilize the scene, their computational cost can be high, and they primarily address the background motion at the association level. In contrast, our HomView-MOT framework advances the CMC paradigm by introducing a highly efficient Fast Homography Estimation (FHE) algorithm that uses keyframes and interpolation to reduce computational burden significantly.
III Method
III-A Overview
@@
III-B Fast Homography Estimation for View Projection
As described in Sec. I, when the UAV flight state changes (e.g. hovering, moving up/down, turning left/right) occur, the captured scene background will undergo significant changes, making adjacent frames appear as different view planes over the same scene. To facilitate cross-view learning and IoU association for object IDs, we propose to perform view projection for these two planes, in the form of homography matrix which is extensively applied to various 3D applications [45]. Generally, the homography matrix for every two adjacent frames can be calculated with the Homography Estimation (HE) algorithm. For instance, as shown in Figure 4, we use a shared SuperPoint [46] network to extract keypoints from both frames separately. The distribution of these keypoints is able to represent the view information of the scene [47, 48, 49, 50]. Then, we can estimate the homography matrix based on the matching of keypoints between the two frames by the SuperGlue algorithm [51].
@@ III-B Fast Homography Estimation for View Projection
However, directly performing homography estimation calculation for every two frames is very time-consuming and would greatly reduce tracking efficiency. Therefore, we propose a Fast Homography Estimation (FHE) method, where we only need to directly calculate the homography matrices for a limited number of key frames, while the homography matrices for other adjacent frames can be fast estimated using the ones of the key frames. Specifically, we utilize a uniform sampling process to select the key frames from the input video sequence $I_{t}(t\in[1,T])$, the sampling interval is set to $h$. We only perform a direct calculation of the homography matrices $\{H_{t,t+h},H_{t+h,t+2h},$...\} between the sampled frames, where $H_{t,t+h}$ indicates the homography matrix from the $t+h$-th to $t$-th frame. For any two frames, based on the following Preliminary of Homography Estimation, we prove that their homography matrix can be reasonably estimated through the derivation of homography matrix of their adjacent key frames.
III-C Derivation of Homography Estimation in Adjacent Frames
When the sampling interval $h$ is sufficiently small, the UAV can be approximated as moving in a nearly linear fashion within this interval. Taking the sampling frames $k_{1},k_{2}(k_{2}=k_{1}+h)$ and any non-sampling frame $t(k_{1}<t<k_{2})$ as an example, for any two points $(x_{k_{1}},y_{k_{1}})$ and $(x_{t},y_{t})$ on the $t$ plane and $k_{1}$ plane respectively, it is approximately true that
$x_{t}=\alpha_{1}x_{k_{2}},~~y_{t}=\alpha_{2}y_{k_{2}},$ (1)
where $\alpha_{1},\alpha_{2}$ are scaling factors based on the proportion of the overall displacement of key points between the two frames to the overall displacement of key points between the adjacent key frames $k_{1},k_{2}$. Formally, let $M_{t,k_{1}}=\{(p_{t}^{1},p_{k_{1}}^{1}),(p_{t}^{2},p_{k_{1}}^{2}),...,(p_{t}^{z},p_{k_{1}}^{z}),...,(p_{t}^{Z_{t,k_{1}}},p_{k_{1}}^{Z_{t,k_{1}}})\}$ denote the matched keypoints between $t$ plane and $k_{1}$ plane, obtained by the SuperPoint + SuperGlue algorithm in Section 3.2 in the main paper. $(p_{t}^{z},p_{k_{1}}^{z})$ indicate the $z$-th matched keypoints in $t$ plane and $k_{1}$ plane respectively. $Z_{t,k_{1}}$ indicates the maximum number of matched point pairs. $p^{z}_{k_{1}}=(x^{z}_{k_{1}},y^{z}_{k_{1}})$, where $x^{z}_{k_{1}},y^{z}_{k_{1}}$ indicate the $x$ and $y$ coordinate of $p^{z}_{k_{1}}$ respectively.
Then, $\alpha_{1},\alpha_{2}$ are scaling factors based on the proportion of the overall displacement of key points between $t$ and $k_{1}$ plane to the overall displacement of key points between the adjacent key frames $k_{1},k_{2}$. Formally,
$\alpha_{1}$ $=\frac{\frac{1}{Z_{t,k_{1}}}\sum_{z=1}^{Z_{t,k_{1}}}D(x_{t}^{z},x_{k_{1}}^{z})}{\frac{1}{Z_{k_{2},k_{1}}}\sum_{z=1}^{Z_{k_{2},k_{1}}}D(x_{k_{2}}^{z},x_{k_{1}}^{z})},$ (2)
$\alpha_{2}$ $=\frac{\frac{1}{Z_{t,k_{1}}}\sum_{z=1}^{Z_{t,k_{1}}}D(y_{t}^{z},y_{k_{1}}^{z})}{\frac{1}{Z_{k_{2},k_{1}}}\sum_{z=1}^{Z_{k_{2},k_{1}}}D(y_{k_{2}}^{z},y_{k_{1}}^{z})},$
where $D(\cdot)$ indicates the L2 distance.
Finally, $\alpha_{t,k_{1}}$ can be calculated by Eq. 9 and Eq. 2. According to the homography property [52], the $t$ plane and $k_{1}$ plane represent different views of the same scene, so we have
$x_{t}=H_{t,k_{1}}x_{k_{1}},~~y_{t}=H_{t,k_{1}}y_{k_{1}}.$ (3)
@@ III-C Derivation of Homography Estimation in Adjacent Frames
From the derivation of homography estimation [52, 45] algorithm, we know that
$A_{t,k_{1}}H_{t,k_{1}}=0,$ (4)
where
$A_{t,k_{1}}=\left[\begin{aligned} &~-x_{k_{1}}~-y_{k_{1}}~~-1~~~~0~~~~~~~~0~~~~~~~~0~~~x_{t}x_{k_{1}}~~x_{t}y_{k_{1}}~~x_{t}~\\ &~~~~0~~~~~~~0~~~~~~~~0~~-x_{k_{1}}~~-y_{k_{1}}~~-1~~y_{t}x_{k_{1}}~~y_{t}y_{k_{1}}~~~y_{t}~\end{aligned}\right].$
Similarly, we have
$A_{k_{2},k_{1}}H_{k_{2},k_{1}}=0,$ (6)
where
$A_{k_{2},k_{1}}=\left[\begin{aligned} &~-x_{k_{1}}~-y_{k_{1}}~~-1~~~~0~~~~~~~~0~~~~~~~~0~~~x_{k_{2}}x_{k_{1}}~~x_{k_{2}}y_{k_{1}}~~x_{k_{2}}~\\ &~~~~0~~~~~~~0~~~~~~~~0~~-x_{k_{1}}~~-y_{k_{1}}~~-1~~y_{k_{2}}x_{k_{1}}~~y_{k_{2}}y_{k_{1}}~~~y_{k_{2}}~\end{aligned}\right].$
Through the Eq. 5, Eq. 7 and Eq. 1, we can derive that
$A_{t,k_{1}}=\alpha_{t,k_{1}}A_{k_{2},k_{1}},$ (8)
where
$\alpha_{t,k_{1}}=\left[\begin{aligned} &~1~~1~~1~~1~~1~~1~~\alpha_{1}~~\alpha_{1}~~\alpha_{1}~\\ &~1~~1~~1~~1~~1~~1~~\alpha_{2}~~\alpha_{2}~~\alpha_{2}~\end{aligned}\right].$ (9)
Next, through Eq. 8, Eq. 4 and Eq. 6, we can get,
$A_{t,k_{1}}H_{t,k_{1}}=0=\alpha_{t,k_{1}}A_{k_{2},k_{1}}H_{k_{2},k_{1}},$ (10)
finally, we can derive that,
$H_{t,k_{1}}=\alpha_{t,k_{1}}H_{k_{2},k_{1}}.$ (11)
III-D Efficient Video-Level Homography Transformation
Specifically, as shown in Figure 5, We select three highly representative frames $I_{t_{1}},I_{t_{2}},I_{t_{3}}$ to illustrate the computation process of the homography matrix between any two frames in the input video, where $k_{1}<t_{1}<t_{2}<k_{2}<t_{3}<k_{3}$, and $k_{1},k_{2},k_{3}$ are the sampled frames. The core calculations are as follows:
@@ III-D Efficient Video-Level Homography Transformation
    • From a non-sampled frame to any frame in the previous sampling interval: $H_{t_{2},t_{1}}=H_{t_{2},k_{2}}H_{k_{2},t_{1}}$ and $H_{t_{2},k_{1}}=H_{t_{2},k_{2}}H_{k_{2},k_{1}}$. The calculation of $H_{t_{3},k_{2}}$ is similar to the previous criterion.
The aforementioned computation process involves only interpolation calculations and matrix multiplication, which ensures high computational efficiency in practical implementation. Through the derivations described above, we can rapidly and cost-effectively calculate the homography matrix between any two frames. Specifically, the parameter $h$ directly controls the balance: 1) Higher Accuracy (Smaller $h$), a smaller $h$ means more frequent direct calculations, leading to higher accuracy but increased computational cost; 2) Higher Efficiency (Larger $h$), A larger $h$ significantly reduces computational load, improving speed (FPS) while accepting a minor approximation.
III-E View-Centric ID Learning
@@
${\rm HSA}_{t}$ $={\rm Softmax}(\frac{k(F_{t}^{ID})(H_{t,t+1}q(S_{t})^{\top})}{\sqrt{C}})v(F_{t}^{ID}),$ (12)
${\rm HSA}_{t+1}$ $={\rm Softmax}(\frac{k(F_{t+1}^{ID})(H_{t+1,t}q(S_{t+1})^{\top})}{\sqrt{C}})v(F_{t+1}^{ID}),$
where $q(\cdot),k(\cdot),v(\cdot)$ are the linear projections respectively. $F_{t}^{ID},F_{t+1}^{ID}$ are the object ID features in frame $t,t+1$, respectively. $S_{t},S_{t+1}$ are the View-Centric Slots in frame $t,t+1$, respectively. $\top$ is the matrix transposition operator. $C$ is the feature dimension. Then we correlate the cross-frame views by,
$S^{\prime}={\rm Softmax}(\frac{S_{t}S_{t+1}^{\top}}{\sqrt{C}})S_{t+1}$ (13)
@@
III-F Homography Matching Filter